# Compositional Difference-in-Differences for categorical outcomes

### v1

@@ 4.1. Impact of early voting law on distributions of voters’ support across parties
Early voting allows registered voters to cast their ballots in person or by mail before Election Day. We evaluate the effect of this policy on voter decisions using the compositional difference-in-differences (CoDiD) framework. A credible analysis requires comparing states with similar political, social, and demographic characteristics so that observed differences in voter participation can be attributed to early voting rather than to preexisting trends. The treatment group includes Maryland and New Jersey, which introduced early voting before 2008, while Pennsylvania and New York serve as the control group. These states are geographically close, share comparable demographic and political profiles, and consistently supported Democratic presidential candidates from 1992 to 2004, indicating similar partisan alignment and voting behavior during the pre-treatment period. Table 2 reports their racial and ethnic composition around 2007 to further contextualize the comparison.
@@
This table shows that the two groups have almost similar racial compositions. Combined with their geographic proximity and political alignment over the study period, this makes them well-suited for causal analysis. The outcome of interest is the voter’s support across three categories, Democrats, Republicans, and Others. We consider that from 1992 elections to 2008. These data are publicly available and consistently reported across states. We start by plotting the log-counts to assess the plausibility of parallel growths.
@@
Our method appears to be the most appropriate in this context, as it relies on the visual assessment of parallel trends between the treatment and control groups. The observed similarity in pre-treatment trajectories provides reassuring evidence that the identifying assumptions are reasonable, thereby strengthening the credibility of the causal analysis.
Empirical findings We begin by examining how the treatment affected the total number of votes cast for each party. Table 3 reports the estimated Growth Treatment Effects (GTT), that is, the proportional increase in votes due to the treatment, along with 95% confidence intervals obtained via the bootstrap procedure described in appendix C.
@@
Overall, the policy increased turnout by around $4.4\%$, demonstrating that early voting can mobilize more voters. This finding is consistent with the existing literature on the effects of early voting. All three categories experienced a statistically significant increase in votes, although the magnitude of the effect varies considerably. The Democratic Party saw a modest 5.5% increase, while the Republican Party experienced a smaller rise of about 2.1%. In contrast, support for Other parties, which includes third-party and independent candidates, grew by nearly 37.5%, representing a much larger relative gain. Importantly, because “Other” parties started from a much smaller base, even a modest absolute increase translates into a large percentage change. Focusing on the two major parties, the Democratic Party benefited more from the policy than the Republican Party. This aligns with the results in Berry et al. (2025), which also found that such policies tend to favor Democrats over Republicans. However, we exercise caution in interpreting these results, as the states we selected were already Democratic-leaning during this period.
@@
Table 4 presents the estimated CTT for the 2008 election. Each entry represents the post-treatment probability of selecting a given party, under a counterfactual scenario where, before treatment, all parties were equally likely (i.e., each had a baseline probability of $1/3$). The results show a clear reallocation of voter preference. After the treatment: the probability of voting for “Other” rises to 38.7%, up from the baseline of 33.3%. The probability of voting Republican falls to 30.0%. The probability of voting Democratic also declines slightly, to 31.4%, but remains higher than that of Republicans. Thus, while both major parties lose relative appeal, Republicans lose more ground than Democrats, and third-party candidates gain the most in relative terms. This suggests the treatment not only increased overall turnout but also reshaped the internal composition of voter choice, pulling support more toward alternative options. This finding is interesting because most prior work focuses exclusively on the two major parties. Our analysis reveals that policy interventions or informational treatments can meaningfully affect the electoral relevance of minor parties, a dimension often overlooked in the literature. It is worth emphasizing that this shift is purely relative: the overall ranking of parties by actual vote share in the treated states remains unchanged (Democrats first, Republicans second, Others third). Hence, the treatment does not overturn the existing political hierarchy, but it does narrow the gap between major and minor parties in terms of voter consideration.
@@ 4.2. Impact of RGGI on electricity generation composition
In this section, we study how the Regional Greenhouse Gas Initiative (RGGI) affected the mix of electricity generation in participating states. RGGI, launched in 2009, is a cooperative, market-based program among Connecticut, Delaware, Maine, Maryland, Massachusetts, New Hampshire, New York, Rhode Island, and Vermont. Its goal is to cap and reduce carbon emissions from the power sector. The program works by setting a regional cap on total carbon dioxide emissions from power plants and issuing a limited number of tradable allowances. Each allowance gives a power plant the right to emit one short ton of carbon dioxide. States distribute these allowances through quarterly auctions. The revenue from these auctions is reinvested in energy efficiency, renewable energy, and other programs that benefit consumers. Evaluating the impact of the Regional Greenhouse Gas Initiative (RGGI) on electricity generation is an interesting causal question because it reveals whether the policy effectively shifts energy production toward cleaner sources and reduces carbon emissions. we analyze U.S. electricity generation data, which is publicly available on the U.S. Energy Information Administration’s website (https://www.eia.gov/electricity/data/state/). This dataset provides annual electricity generation figures in megawatts by energy source for each state. We consider the following four categories of energy sources: gas (combining natural gas and other gases), coal and oil (combining coal and petroleum), nuclear (corresponding solely to the nuclear category), and renewables (combining hydroelectric, conventional, solar thermal and photovoltaic, geothermal, wind, and wood and wood-derived fuel). For each state and year, we observe the number of megawatts generated by different sources, and we aggregate these into two groups: the treated group and the control group. In constructing the control group, we exclude states that either joined RGGI, are strongly affected by leakage through electricity market integration, or implemented their own carbon pricing policies. In particular, Pennsylvania, West Virginia, Virginia, and Ohio are excluded due to their geographic proximity and electricity market linkages with RGGI states, which expose them to leakage and anticipatory effects. In fact, electricity markets in the U.S. Northeast and Mid-Atlantic are highly integrated (PJM, NYISO, ISO-NE), including these states in the control group risks contamination. California, Washington, and Oregon are also excluded, since those states adopted cap-and-trade or cap-and-invest programs that constitute carbon pricing mechanisms similar in nature to RGGI. Including them in the control pool would compromise the validity of the counterfactual by mixing treated and untreated units.
@@
The program was organized into three-year compliance periods, with the first spanning January 1, 2009, to December 31, 2011. We focus our analysis on this initial period because the parallel growth assumption becomes progressively less plausible over longer time horizons. Here, we also assess the validity of the parallel growths assumption by plotting both the raw quantities and their logarithmic transformations over the study period (2000-2011). As shown in Figure 10, the log-quantity trajectories for the treated and control groups move approximately in parallel during the pre-treatment period, supporting the plausibility of the assumption.
@@
We examine the impact of RGGI over the first three years of implementation (2009–2011), corresponding to the program’s initial compliance period. As shown in the dynamic estimates below, the policy induced a statistically significant and growing reduction in total electricity generation (see figure11).
@@
While total output fell, the composition of generation shifted, as we can see in table 5.
@@ 5. Conclusion
This paper introduces compositional difference-in-differences (CoDiD), a causal inference framework for analyzing how treatments and policies affect entire categorical outcome quantities and distributions. It is especially suited to settings with discrete, unordered outcomes, such as employment status (employed, unemployed, out of the labor force), voting choice, or health categories, where the policy-relevant question is not whether a single average has changed, but how the composition of all categories has been reconfigured. CoDiD addresses a dual objective: it quantifies both absolute changes in total counts and relative shifts in shares (i.e., the probability mass function). This allows researchers and policymakers to identify which categories expanded or contracted, by how much, and in what combination of scale and composition. The framework is built on a parallel growth assumption, a multiplicative analog of standard parallel trends, which posits that, in the absence of treatment, the ratios of category sizes evolve proportionally over time. This assumption has a natural interpretation in random utility models: it corresponds to stable relative utility differences across categories, so that estimated effects map directly to shifts in underlying preferences. CoDiD is designed for practical use and naturally extends to richer settings, including relaxations of the identifying assumption, staggered treatment adoption, and a synthetic DiD analog for richer data contexts. A key direction for future work is the incorporation of continuous covariates, enabling adjustment for confounders and the estimation of heterogeneous treatment effects across subpopulations.
@@ A.1. Proof of Theorem 1
Since all vectors are strictly positive (by Assumption 1), the logarithm is defined component-wise, and we can apply the component-wise exponential function $\exp\cdot$ to both sides of (1). Because $\exp\cdot$ and $\log\cdot$ are inverses on $\mathbb{R}_{>0}^{p}$, we obtain:
@@ A.2. Proof of proposition 1
We consider the untreated counterfactual quantities $q_{g,t}^{N}(c_{k})>0$ for groups $g\in{0,1}$, times $t\in{0,1}$. Pick a baseline category $c_{p}$. Let $S_{g,t}^{N}=\sum_{k=1}^{p}q_{g,t}^{N}(c_{k})$ be the counterfactual total for all category for group $g$ at time $t$. We know that $\pi(c_{k})=\frac{q_{g,t}^{N}(c_{k})}{S_{g,t}^{N}}$. The log-odds for each category can be written as :
@@
A.3. Proof of Proposition 2
We aim to show that the parallel growths assumption in compositional form is equivalent to a parallel trends condition in log-ratio coordinates. Specifically, we prove the equivalence of the following two statements:
$(i)\pi_{1,1}^{N}\ominus\pi_{0,1}^{N}=\pi_{1,0}^{N}\ominus\pi_{0,0}^{N}$
$(ii)\textit{ for all }k=1,\dots,p-1,\log\left(\frac{\pi_{1,1}^{N}(c_{k})}{\pi_{1,1}^{N}(c_{p})}\right)-\log\left(\frac{\pi_{0,1}^{N}(c_{k})}{\pi_{0,1}^{N}(c_{p})}\right)=\log\left(\frac{\pi_{1,0}^{N}(c_{k})}{\pi_{1,0}^{N}(c_{p})}\right)-\log\left(\frac{\pi_{0,0}^{N}(c_{k})}{\pi_{0,0}^{N}(c_{p})}\right).$
@@ A.3. Proof of Proposition 2
We know that $\ell:\text{int}(\mathcal{S}^{p-1})\to\mathbb{R}^{p-1}$ defined by:
@@
This map is a bijection: for any $x=(x_{1},\dots,x_{p-1})\in\mathbb{R}^{p-1}$, define $r_{j}=e^{x_{j}}$ for $j=1,\dots,p-1$, and set
$\pi(c_{j})=\frac{r_{j}}{1+\sum_{g=1}^{p-1}r_{g}},\quad j=1,\dots,p-1,\qquad\pi(c_{p})=\frac{1}{1+\sum_{g=1}^{p-1}r_{g}}.$
A key property of that transformation is that it linearizes the Aitchison geometry of the simplex. In particular, for any $\pi_{1},\pi_{2}\in\text{int}(\mathcal{S}^{p-1})$,

### v2

@@ 4.1. Impact of early voting law on distributions of voters’ support across parties
Early voting allows registered voters to cast their ballots in person or by mail before Election Day. I evaluate the effect of this policy on voter decisions using the compositional difference-in-differences (CoDiD) framework. A credible analysis requires comparing states with similar political, social, and demographic characteristics so that observed differences in voter participation can be attributed to early voting rather than to preexisting trends. The treatment group includes Maryland and New Jersey, which introduced early voting before 2008, while Pennsylvania and New York serve as the control group. These states are geographically close, share comparable demographic and political profiles, and consistently supported Democratic presidential candidates from 1992 to 2004, indicating similar partisan alignment and voting behavior during the pre-treatment period. Table 2 reports their racial and ethnic composition around 2007 to further contextualize the comparison.
@@
This table shows that the two groups have almost similar racial compositions. Combined with their geographic proximity and political alignment over the study period, this makes them well-suited for causal analysis. The outcome of interest is the voter’s support across three categories, Democrats, Republicans, and Others. I consider that from 1992 elections to 2008. These data are publicly available and consistently reported across states. I start by plotting the log-counts to assess the plausibility of parallel growths.
@@
My method appears to be the most appropriate in this context, as it relies on the visual assessment of parallel trends between the treatment and control groups. The observed similarity in pre-treatment trajectories provides reassuring evidence that the identifying assumptions are reasonable, thereby strengthening the credibility of the causal analysis.
Empirical findings I begin by examining how the treatment affected the total number of votes cast for each party. Table 3 reports the estimated Growth Treatment Effects (GTT), that is, the proportional increase in votes due to the treatment, along with 95% confidence intervals obtained via the bootstrap procedure described in appendix C.
@@
Overall, the policy increased turnout by around $4.4\%$, demonstrating that early voting can mobilize more voters. This finding is consistent with the existing literature on the effects of early voting. All three categories experienced a statistically significant increase in votes, although the magnitude of the effect varies considerably. The Democratic Party saw a modest 5.5% increase, while the Republican Party experienced a smaller rise of about 2.1%. In contrast, support for Other parties, which includes third-party and independent candidates, grew by nearly 37.5%, representing a much larger relative gain. Importantly, because “Other” parties started from a much smaller base, even a modest absolute increase translates into a large percentage change. Focusing on the two major parties, the Democratic Party benefited more from the policy than the Republican Party. This aligns with the results in Berry et al. (2025), which also found that such policies tend to favor Democrats over Republicans. However, I exercise caution in interpreting these results, as the states I selected were already Democratic-leaning during this period.
@@
Table 4 presents the estimated CTT for the 2008 election. Each entry represents the post-treatment probability of selecting a given party, under a counterfactual scenario where, before treatment, all parties were equally likely (i.e., each had a baseline probability of $1/3$). The results show a clear reallocation of voter preference. After the treatment: the probability of voting for “Other” rises to 38.7%, up from the baseline of 33.3%. The probability of voting Republican falls to 30.0%. The probability of voting Democratic also declines slightly, to 31.4%, but remains higher than that of Republicans. Thus, while both major parties lose relative appeal, Republicans lose more ground than Democrats, and third-party candidates gain the most in relative terms. This suggests the treatment not only increased overall turnout but also reshaped the internal composition of voter choice, pulling support more toward alternative options. This finding is interesting because most prior work focuses exclusively on the two major parties. My analysis reveals that policy interventions can meaningfully affect the electoral relevance of minor parties, a dimension often overlooked in the literature. It is worth emphasizing that this shift is purely relative: the overall ranking of parties by actual vote share in the treated states remains unchanged (Democrats first, Republicans second, Others third). Hence, the treatment does not overturn the existing political hierarchy, but it does narrow the gap between major and minor parties in terms of voter consideration.
@@ 4.2. Impact of RGGI on electricity generation composition
In this section, I study how the Regional Greenhouse Gas Initiative (RGGI) affected the mix of electricity generation in participating states. RGGI, launched in 2009, is a cooperative, market-based program among Connecticut, Delaware, Maine, Maryland, Massachusetts, New Hampshire, New York, Rhode Island, and Vermont. Its goal is to cap and reduce carbon emissions from the power sector. The program works by setting a regional cap on total carbon dioxide emissions from power plants and issuing a limited number of tradable allowances. Each allowance gives a power plant the right to emit one short ton of carbon dioxide. States distribute these allowances through quarterly auctions. The revenue from these auctions is reinvested in energy efficiency, renewable energy, and other programs that benefit consumers. Evaluating the impact of the Regional Greenhouse Gas Initiative (RGGI) on electricity generation is an interesting causal question because it reveals whether the policy effectively shifts energy production toward cleaner sources and reduces carbon emissions. I analyze U.S. electricity generation data, which is publicly available on the U.S. Energy Information Administration’s website (https://www.eia.gov/electricity/data/state/). This dataset provides annual electricity generation figures in megawatts by energy source for each state. I consider the following four categories of energy sources: gas (combining natural gas and other gases), coal and oil (combining coal and petroleum), nuclear (corresponding solely to the nuclear category), and renewables (combining hydroelectric, conventional, solar thermal and photovoltaic, geothermal, wind, and wood and wood-derived fuel). For each state and year, I observe the number of megawatts generated by different sources, and I aggregate these into two groups: the treated group and the control group. In constructing the control group, I exclude states that either joined RGGI, are strongly affected by leakage through electricity market integration, or implemented their own carbon pricing policies. In particular, Pennsylvania, West Virginia, Virginia, and Ohio are excluded due to their geographic proximity and electricity market linkages with RGGI states, which expose them to leakage and anticipatory effects. In fact, electricity markets in the U.S. Northeast and Mid-Atlantic are highly integrated (PJM, NYISO, ISO-NE), including these states in the control group risks contamination. California, Washington, and Oregon are also excluded, since those states adopted cap-and-trade or cap-and-invest programs that constitute carbon pricing mechanisms similar in nature to RGGI. Including them in the control pool would compromise the validity of the counterfactual by mixing treated and untreated units.
@@
The program was organized into three-year compliance periods, with the first spanning January 1, 2009, to December 31, 2011. I focus my analysis on this initial period because the parallel growth assumption becomes progressively less plausible over longer time horizons. Here, I also assess the validity of the parallel growths assumption by plotting both the raw quantities and their logarithmic transformations over the study period (2000-2011). As shown in Figure 10, the log-quantity trajectories for the treated and control groups move approximately in parallel during the pre-treatment period, supporting the plausibility of the assumption.
@@
I examine the impact of RGGI over the first three years of implementation (2009–2011), corresponding to the program’s initial compliance period. As shown in the dynamic estimates below, the policy induced a statistically significant and growing reduction in total electricity generation (see figure11).
@@
While total output fell, the composition of generation shifted, as I can see in table 5.
@@ 5. Conclusion
This paper introduces compositional difference-in-differences (CoDiD), a causal inference framework for analyzing how treatments and policies affect entire categorical outcome quantities and distributions. It is especially suited to settings with discrete, unordered outcomes, such as employment status (employed, unemployed, out of the labor force), voting choice, or health categories, where the policy-relevant question is not whether a single average has changed, but how the composition of all categories has been reconfigured. CoDiD addresses a dual objective: it quantifies both absolute changes in total counts and relative shifts in shares (i.e., the probability mass function). This allows researchers and policymakers to identify which categories expanded or contracted, by how much, and in what combination of scale and composition. The framework is built on a parallel growths assumption, a multiplicative analog of standard parallel trends, which posits that, in the absence of treatment, the ratios of category sizes evolve proportionally over time. This assumption has a natural interpretation in random utility models: it corresponds to stable relative utility differences across categories, so that estimated effects map directly to shifts in underlying preferences. CoDiD is designed for practical use and naturally extends to richer settings, including relaxations of the identifying assumption, staggered treatment adoption, and a synthetic DiD analog for richer data contexts. A key direction for future work is the incorporation of continuous covariates, enabling adjustment for confounders and the estimation of heterogeneous treatment effects across subpopulations.
@@ A.1. Proof of Theorem 1
Since all vectors are strictly positive (by Assumption 1), the logarithm is defined component-wise, and I can apply the component-wise exponential function $\exp\cdot$ to both sides of (1). Because $\exp\cdot$ and $\log\cdot$ are inverses on $\mathbb{R}_{>0}^{p}$, I obtain:
@@ A.2. Proof of proposition 1
we have:
$U^{N}_{gt}(c_{k})=\mu^{N}_{gt}(c_{k})-\log\left(\sum_{k}e^{\mu^{N}_{gt}(c_{k})}\right)+\log\left(S^{N}_{gt}\right)+\varepsilon_{gt}(c_{k})$
In the standard multinomial logit RUM:
$\pi_{gt}(c_{k})=\Pr\Big(U^{N}_{gt}(c_{k})\geq U^{N}_{gt}(c_{j})\text{ for all }j\neq k\Big).$
Since $\varepsilon_{gt}(c_{k})\sim\text{i.i.d. Gumbel}$, the choice probability has the closed-form logit formula:
$\pi_{gt}(c_{k})=\frac{\exp\Big(\mu^{N}_{gt}(c_{k})+\log(S^{N}_{gt})-\log(\sum_{j}e^{\mu^{N}_{gt}(c_{j})})\Big)}{\sum_{j}\exp\Big(\mu^{N}_{gt}(c_{j})+\log(S^{N}_{gt})-\log(\sum_{l}e^{\mu^{N}_{gt}(c_{l})})\Big)}.$
This simplifies to:
$\pi_{gt}(c_{k})=\frac{e^{\mu^{N}_{gt}(c_{k})}\cdot\frac{S^{N}_{gt}}{\sum_{l}e^{\mu^{N}_{gt}(c_{l})}}}{\sum_{j}e^{\mu^{N}_{gt}(c_{j})}\cdot\frac{S^{N}_{gt}}{\sum_{l}e^{\mu^{N}_{gt}(c_{l})}}}=\frac{e^{\mu^{N}_{gt}(c_{k})}}{\sum_{j}e^{\mu^{N}_{gt}(c_{j})}}.$
The $S^{N}_{gt}$ term cancels out in the probability, which shows that the relative shares depend only on the relative utilities, while $S^{N}_{gt}$ scales the absolute number of choices. The number of individuals choosing $c_{k}$ is:
$q_{gt}^{N}(c_{k})=\pi_{gt}(c_{k})\cdot S^{N}_{gt}=\frac{e^{\mu^{N}_{gt}(c_{k})}}{\sum_{j}e^{\mu^{N}_{gt}(c_{j})}}\cdot S^{N}_{gt}.$
This leads to:
$\log q_{gt}^{N}(c_{k})=\mu_{gt}^{N}(c_{k})-\log\left(\sum_{k=1}^{p}e^{\mu_{gt}^{N}(c_{k})}\right)+\log(S_{gt}^{N}).$
Now, we have :
$\mathbb{E}[U_{g1}(c_{k})]-\mathbb{E}[U_{g0}(c_{k})]$ $=$ $\mu^{N}_{g1}(c_{k})-\log(\sum_{k}e^{\mu^{N}_{g1}(c_{k})})+\log\left(S^{N}_{g,1}\right)+\gamma$
$-\mu^{N}_{g0}(c_{k})-\log(\sum_{k}e^{\mu^{N}_{g0}(c_{k})})+\log\left(S^{N}_{g0}\right)-\gamma$
$=$ $\log q_{g1}^{N}(c_{k})-\log q_{g1}^{N}(c_{k})$
From there, we have that assumption 2 is equivalent to parallel trends in expected utilities:
$\mathbb{E}[U^{N}_{11}(c_{k})]-\mathbb{E}[U^{N}_{10}(c_{k})]=\mathbb{E}[U^{N}_{01}(c_{k})]-\mathbb{E}[U^{N}_{00}(c_{k})],\quad\forall c_{k}\in\bar{\mathcal{Y}}.$
A.3. Proof of Proposition 2
Pick a baseline category $c_{p}$. The log-odds for each category can be written as :
@@ A.3. Proof of Proposition 2
From part 1, I also can derive that:
$\big\{\log(q_{1,1}^{N}(c_{k}))-\log(q_{1,1}^{N}(c_{l}))\big\}-\big\{\log(q_{1,0}^{N}(c_{k}))-\log(q_{1,0}^{N}(c_{l}))\big\}=$
$\big\{\log(q_{0,1}^{N}(c_{l}))-\log(q_{0,1}^{N}(c_{l}))\big\}-\big\{\log(q_{0,0}^{N}(c_{k}))-\log(q_{0,0}^{N}(c_{l})).\big\}$
is equivalent to
$\Big\{\mathbb{E}[U^{N}_{11}(c_{k})]-\mathbb{E}[U^{N}_{11}(c_{l})]\Big\}-\Big\{\mathbb{E}[U^{N}_{10}(c_{k})]-\mathbb{E}[U^{N}_{10}(c_{l})]\Big\}=$
$\Big\{\mathbb{E}[U^{N}_{01}(c_{k})]-\mathbb{E}[U^{N}_{01}(c_{l})]\Big\}-\Big\{\mathbb{E}[U^{N}_{00}(c_{k})]-\mathbb{E}[U^{N}_{00}(c_{l})]\Big\},\quad\forall\,c_{k},c_{l}\in\bar{\mathcal{Y}}.$ (14)
A.4. Proof of Proposition 3
I aim to show that the parallel growths assumption in compositional form is equivalent to a parallel trends condition in log-odds. Specifically, I prove the equivalence of the following two statements:
$(i)\quad\pi_{1,1}^{N}\ominus\pi_{0,1}^{N}=\pi_{1,0}^{N}\ominus\pi_{0,0}^{N}$
$(ii)\quad\textit{ for all }k=1,\dots,p-1,\log\left(\frac{\pi_{1,1}^{N}(c_{k})}{\pi_{1,1}^{N}(c_{p})}\right)-\log\left(\frac{\pi_{0,1}^{N}(c_{k})}{\pi_{0,1}^{N}(c_{p})}\right)=\log\left(\frac{\pi_{1,0}^{N}(c_{k})}{\pi_{1,0}^{N}(c_{p})}\right)-\log\left(\frac{\pi_{0,0}^{N}(c_{k})}{\pi_{0,0}^{N}(c_{p})}\right).$
@@ A.4. Proof of Proposition 3
I know that $\ell:\text{int}(\mathcal{S}^{p-1})\to\mathbb{R}^{p-1}$ defined by:
@@
I already know that this map is a bijection. A key property of that transformation is that it linearizes the Aitchison geometry of the simplex. In particular, for any $\pi_{1},\pi_{2}\in\text{int}(\mathcal{S}^{p-1})$,